A solution contains 0.236 M potassium hypochlorite and 6.74×10-2 M hypochlorous acid.

The pH of this solution is
Answer #1

Dear Student,

Please see to the solution and If there are any doubts or need further explanation,please feel free to comment here and may upvote if you understand the solution.

The answer is pH=8.00

Given a buffer solution of weak acid and it's conjugate base.

Potassium hypochlorite (KOCl) will give K+ and OCl- ions in aqueous solution.

K+ will not have any effect on pH of solution and   OCl- ion reacts with water and forms HClO (HYPOCHLOROUS ACID).

From Henderson Equation of a Buffer consisting weak acid and it's conjugate base We have formulae to determine pH ofsolution

pH=pKa+log{ (Conjugate Base) / (Weak acid) }

pKa of HClO is 7.46 which is a standard value

pH= 7.46 + log{ (0.236)/(0.0674) } (6.74*10^(-2)=0.0674)

pH= 7.46+0.54

pH= 8.00

The answer is pH=8.00